4. A golf ball has a diameter of 4.27 cm.

(a) If a golf player hits the ball into a stream, what upthrust does it experience when it is completely submerged? (Assume density of water = 1000 kg m−3.)

(b) If the mass of the ball is 45 g, what is the resultant force on it when underwater?

(c) Referring to Newton’s laws of motion, explain what will happen to the submerged golf ball.


1
Expert's answer
2020-12-24T11:42:14-0500

Answer

(a)

Upthrust in ball can be given as

"F_t=V\rho g\\=\frac{4}{3}\pi r^3\rho g\\=\frac{4}{3}\pi (4.27\times10^{-2}/2)^3(1000) (9.8) \\F_t=3.99N"


b) resultant force

"F_R=F_t-mg\\=3.99-(0.045\times9.8) =3.5N"



(C) according to law of newton

"F_R=ma\\a=\frac{3.5}{0.045}=77.7m/s^2"


Ball will move in upward direction with acceleration of "77.7m\/s^2"